# Axi-majoron for almost everything

### v2

@@ 3.3 Reheating after inflation
For $\phi$ well below the Planck scale, the scalar potential is approximated as $U\simeq{\tilde{V}}_{S}$. It can be expressed as a familiar form given in Eq. (2.16). For simplicity, we consider the case of $\lambda_{\phi}^{\prime}=\lambda_{\phi}$ (i.e., $\zeta_{h\phi}=0\ \Leftrightarrow\lambda_{h\phi}=0$). Then, the mass of $\phi$-quanta for a given background value is given by
@@
For $\lambda_{h\phi}=0$ the production of the SM Higgs from either preheating processes or decays of $\phi$ particles is negligible. For sufficiently small $y_{N}$ which will be discussed in the subsequent discussion, the production of RHNs via preheating is also not efficient [56]. Hence, initially there can be population of $\phi$-particles and axi-majorons ($a_{\phi}$) only  [^16] There are thermal particles with a temperature $T\sim H_{\rm e}/(2\pi)$ at the end of inflation [57]. We ignore this contribution which does not affect our argument. . As shown in Ref. [58], the self-preheating effect on $\phi$-particles is inefficient. The population of $a_{\phi}$ is only from the perturbative decay of $\phi$-particles through the kinetic term of $a_{\phi}$. Moreover, if $\Phi$ has non-zero angular momentum (i.e., a phase rotation), the preheating effect will be suppressed since $\phi$ does not pass through the origin during its oscillations and the violation of adiabaticity which is necessary for an efficient preheating becomes weak as a result. However, there might be a possibility of efficient energy transfer via axi-majoron’s anomalous couplings to non-Abelian gauge fields due to large velocity of axi-majoron after inflation [59]. Since the velocity of axi-majoron is periodically changing as shown in Fig. 3, in regard of the polarization state of gauge field which is amplified the situation is similar to the case of parametric resonance in a conformal scalar theory [58]. The details will be investigated in other work. In any case, as long as the eventual decay of $\phi$-particle is the dominant source of the energy of the universe and the interaction of $\phi$ to RHNs is small enough to avoid the possibility of thermal-trapping, our argument is barely affected. Hence, in this work we ignore the effect of preheating.
In our scenario, $U(1)_{\rm PQ}$ symmetry is broken only in the gravity sector, and its effect is suppressed by $\alpha\xi_{\phi}\phi^{2}/M_{\rm P}^{2}$. As a result, $a_{\phi}$ is lighter than $\phi$ for $\phi\lesssim M_{\rm P}$ with $\alpha\xi_{\phi}\ll 1$, and $\phi$ can always decay to $a_{\phi}$s via the kinetic term of $a_{\phi}$. The partial decay rate of the channel is given by
@@
From the seesaw formula for the mass parameters of the left-handed neutrinos [60, 61]  [^17] At low energy below the electroweak scale, $y_{\nu}$ should be replaced to the one obtained from a bi-unitarity transformation associated with physical mass-eigenstates. We ignore such details here, since it does not make changes of the main point of our argument. In this reason, we assumed $y_{\nu}$ is flavor-diagonal for simplicity.,
$m_{\nu_{i}}\approx\frac{y_{\nu_{i}}^{2}v_{h}^{2}}{2m_{N_{i}}}\ .$ (3.57)
Then the constraint in Eq. (3.56) is translated as a constraint on $y_{\nu_{i}}$, which can be expressed as
$y_{\nu_{i}}^{2}$$\approx$ $2\sqrt{b_{i}\lambda_{\phi}}\left(\frac{m_{\nu_{i}}\phi_{0}}{v_{h}^{2}}\right)\ ,$ (3.58)
@@
For $\phi\gtrsim\phi_{\times}$ the universe is driven by a radiation-like energy component. In this case, the decay can happen effectively once $H\approx\Gamma^{\phi}/2$ is satisfied. If the subdominant angular motion is ignored, the expansion rate after inflation can be expressed such as
@@
$\frac{\phi_{*}}{\phi_{\times}}\approx\frac{9\lambda_{\phi}}{32\sqrt{2}\pi}\frac{M_{\rm P}}{\phi_{0}}\simeq\left(\frac{\xi_{\phi}}{10^{2}}\right)^{2}\left(\frac{\phi_{0}^{\rm ref}}{\phi_{0}}\right)\ .$ (3.63)
@@
where ${\tilde{\phi}}$ is understood as the amplitude of oscillation. In this regime, the universe is in a matter-domination era driven by non-relativistic massive $\phi$-particles [^18] There is no kination era driven by the kinetic anergy caused by angular motion of $\Phi$. This is because the phase velocity of $\Phi$, i.e., $\dot{\theta}$ scales as $a^{-1}$ for $\phi\gtrsim\phi_{\times}$ and $a^{-3}$ for $\phi\lesssim\phi_{\times}$ in our scenario. Its energy contribution is always sub-dominant relative to the contribution from the radial mode.. Hence, requiring $H=(2/3)\Gamma^{\phi}$, one finds the oscillation amplitude ${\tilde{\phi}}_{*}$ at the epoch of the decay of $\phi$ as
${\tilde{\phi}}_{*}\approx\frac{\sqrt{6}\lambda_{\phi}}{24\pi}M_{\rm P}\simeq 4.7\times 10^{7}{\rm GeV}\times\xi_{\phi}^{2}\ ,$ (3.66)
@@
${\tilde{H}}_{*}\approx\frac{\sqrt{2}\lambda_{\phi}^{3/2}}{24\pi}\phi_{0}\simeq 3.5\times 10^{-4}{\rm GeV}\times\xi_{\phi}^{3}\left(\frac{\phi_{0}}{\phi_{0}^{\rm ref}}\right)\ .$ (3.67)
@@
$\xi_{\phi}\lesssim 680\left(\frac{0.1}{b_{N}}\right)^{1/5}\left(\frac{\phi_{0}}{\phi_{0}^{\rm ref}}\right)^{2/5}$ (3.70)
where we took $b_{N}\equiv b_{3}\approx b_{2}\gg b_{1}$. In the case of $\phi_{*}<\phi_{\times}$, the constraint on $\xi_{\phi}$ is much weaker since the would-be temperature of RHNs around the epoch of $\phi=\phi_{\times}$ is much lower. Hence, conservatively speaking, as long as Eq. (3.70) is satisfied, the PQ-field can avoid thermal-trap around the origin, and the PQ-symmetry is never restored during and after inflation.
@@
The details of the post-inflation cosmology depends on whether $\phi_{*}$ is larger than $\phi_{\times}$ or not. So, we discuss those cases separately.
@@
$\xi_{\phi}<100\times\left(\frac{\phi_{0}}{\phi_{0}^{\rm ref}}\right)^{1/2}$ (3.71)
Once $\phi$ decays at the epoch of ${\tilde{H}}_{*}$, the universe will be filled with axi-majorons and RHNs as a comparable or sub-dominant component. Depending on interaction rates, they may or may not reach thermal or kinetic equilibrium among themselves before the standard thermal bath is established from decays of RHNs. Initially, the production of $\phi$ via inverse decay is kinematically forbidden. Hence, thermalization before the decay of RHNs would be driven by interactions of axi-majoron kinetic term, $y_{N_{i}}$, and $y_{\nu_{i}}$. The $\phi$-mediating self-interaction cross-section of axi-majorons is given by
@@
It implies that $E_{a,\rm k}>{\tilde{E}}_{a,*}$ if $\xi_{\phi}\lesssim 100\left(\phi_{0}/\phi_{0}^{\rm ref}\right)^{1/2}$. Hence, for $\xi_{\phi}$ satisfying Eq. (3.71), axi-majorons are never thermalized among themselves after their production in the decay of $\phi$.
@@
where $N=\left(\nu_{R},\nu_{R}^{c}\right)^{T}$ is a four-component Majorana field with the flavor index suppressed, and only relevant leading order terms are shown. When the condition in Eq. (3.56) is satisfied for the RHN-channels of the decay of the inflaton, the cross-sections of the various interaction between the axi-majorons and the RHNs are smaller than the one for axi-majoron self-interactions at high energy limit($s\approx m_{\phi,0}$) at least by a factor of about $b_{i}/3$ (see Appendix C). Moreover, the number density of RHNs is also suppressed by ${\rm B}_{i}$. As a result, for $\xi_{\phi}$ satisfying Eq. (3.71), the associated interaction rates can not catch up the expansion rate for $H<H_{\times}\equiv H(\phi_{\times})$.
@@
$\frac{y_{\nu_{i}}^{4}}{y_{N_{i}}^{4}}=\left[\frac{m_{\nu}\phi_{0}}{v_{h}^{2}\sqrt{b_{i}\lambda_{\phi}}}\left(\frac{m_{\nu_{i}}}{m_{\nu}}\right)\right]^{2}\approx\frac{10^{3}}{b_{i}\xi_{\phi}^{2}}\left[\frac{\phi_{0}}{\phi_{0}^{\rm ref}}\frac{m_{\nu_{i}}}{m_{\nu}}\right]^{2}\ ,$ (3.77)
one may think that the interaction $y_{\nu_{i}}$ can cause much more efficient scatterings and annihilations of $N_{i}$s to SM particles. However, the decay rate of a RHN mass eigenstate $N_{i}$ is given by (see Appendix D)
$\Gamma_{N_{i}}\simeq\frac{y_{\nu_{i}}^{2}m_{N_{i}}}{8\pi}=\frac{m_{\nu_{i}}}{4\pi}\left(\frac{m_{N_{i}}}{v_{h}}\right)^{2}=\frac{b_{i}\lambda_{\phi}m_{\nu_{i}}}{4\pi}\left(\frac{\phi_{0}}{v_{h}}\right)^{2}\ ,$ (3.78)
where $m_{N_{i}}\gg m_{h}$ was assumed. Also, if the left-handed neutrino mass-eigenstates have the normal hierarchy [15] for their masses as an example, $m_{\nu_{i}}\approx m_{\nu}$ is required for the corresponding heavy RHN mass-eigenstates which we assume to be $N_{2,3}$. Then, for $\xi_{\phi}$ in the range we are interested in, with $\phi_{0}\sim\phi_{0}^{\rm ref}$, heavy RHN mass-eigenstates decay well before their self-interactions/annihilations involving either axi-majorons or SM particles become relevant. The lightest RHNs can also be decoupled from SM particles before their decay if $m_{\nu_{1}}\ll m_{\nu}$ even in the case of $\xi_{\phi}\ll\mathcal{O}(10^{2})$. Therefore, once produced, axi-majorons and RHNs are expected to evolve independently without thermalization at least until the epoch of RHNs decay as long as $m_{\nu_{1}}\ll m_{\nu_{2}}\approx m_{\nu_{3}}\approx m_{\nu}$.
If it were stable, a RHN mass-eigenstate becomes non-relativistic when the expansion rate becomes
@@
$\frac{\Gamma_{N_{i}}}{{\tilde{H}}_{i,\rm non}}=\frac{3}{\sqrt{2\lambda_{\phi}}}\frac{m_{\nu_{i}}\phi_{0}}{v_{h}^{2}}\simeq\frac{66}{\xi_{\phi}}\left(\frac{m_{\nu_{i}}}{m_{\nu}}\right)\left(\frac{\phi_{0}}{\phi_{0}^{\rm ref}}\right)\ .$ (3.80)
@@
$\frac{\Gamma_{N_{i}}}{{\tilde{H}}_{i,\rm eq}}=\frac{3}{2\sqrt{\lambda_{\phi}}{\rm B}_{i}^{2}}\left(\frac{m_{\nu_{i}}\phi_{0}}{v_{h}^{2}}\right)\approx\frac{51}{{\rm B}_{i}^{2}\xi_{\phi}}\left(\frac{m_{\nu_{i}}}{m_{\nu}}\right)\left(\frac{\phi_{0}}{\phi_{0}^{\rm ref}}\right)\ .$ (3.83)
If one of RHN mass-eigenstates is responsible for eventual recovering the SM thermal bath, it would have to satisfy $\Gamma_{N_{i}}/{\tilde{H}}_{i,\rm eq}<0.1\times\left(\Delta N_{\rm eff}^{\rm obs}/0.5\right)^{3/2}$ with $\Delta N_{\rm eff}^{\rm obs}$ being the observational bound on the extra radiation (see Sectoin 5). Combined with the condition in Eq. (3.71), the branching ratio ${\rm B}_{i}$ is then constrained as
${\rm B}_{i}\gtrsim 0.40\times\left(\frac{0.5}{\Delta N_{\rm eff}^{\rm obs}}\right)^{3/4}\left(\frac{m_{\nu_{i}}}{m_{\nu}}\right)^{1/2}\left(\frac{\phi_{0}}{10^{9}{\rm GeV}}\right)^{1/4}\ ,$ (3.84)
@@
$\frac{m_{\nu_{1}}}{m_{\nu}}<2\times 10^{-4}\times\left(\frac{\Delta N_{\rm eff}^{\rm obs}}{0.5}\right)^{3/2}\left(\frac{\xi_{\phi}}{10}\right)\left(\frac{{\rm B}_{1}}{0.1}\right)^{2}\left(\frac{\phi_{0}}{\phi_{0}^{\rm ref}}\right)\ .$ (3.85)
Then, it is necessary to have $\phi_{0}\sim\phi_{0}^{\rm ref}$ for the observed relic density of dark matter (see Section 5), and heavy states are expected to decay around or before the epoch they become non-relativistic, and never dominate the universe.
If heavy RHN states decay before or around the epoch they become non-relativistic, their decay rate is suppressed by the boosting effect. For simplicity, we may take ${\rm B}_{N}\equiv{\rm B}_{3}\approx{\rm B}_{2}$, $\Gamma_{N}\equiv\Gamma_{N_{3}}\approx\Gamma_{N_{2}}$, $m_{\nu_{N}}\equiv m_{\nu_{3}}\approx m_{\nu_{2}}$, $m_{N}\equiv m_{N_{3}}\approx m_{N_{2}}$. Then, the expansion rate at the epoch of the decays of heavy RHNs is estimated as
$H_{N}=\frac{f_{N}(\xi_{\phi},\phi_{0})}{2}\Gamma_{N}\ ,$ (3.86)
@@
$f_{N}(\xi_{\phi},\phi_{0})\equiv\left[\frac{{\tilde{H}}_{*}}{\Gamma_{N}/2}\left(\frac{m_{N}}{m_{\phi,0}}\right)^{2}\right]^{1/3}\simeq 0.3\xi_{\phi}^{1/3}\left(\frac{\phi_{0}^{\rm ref}}{\phi_{0}}\right)^{1/3}\ ,$ (3.87)
@@
If $\phi_{*}>\phi_{\times}$, i.e., $\xi_{\phi}>100\sqrt{\phi_{0}/\phi_{0}^{\rm ref}}$ (see Eq. (3.63)), $\Phi$ field decays before falling into the region of the true vacuum position. Even in this case, $N_{i}$ can be kept out of kinetic or thermal equilibrium as long as $\xi_{\phi}\sim\mathcal{O}(10^{2})$ with $\phi_{0}\approx\phi_{0}^{\rm ref}$. The expansion rate at the epoch when $N_{i}$ becomes non-relativistic is found as
@@
$\frac{\Gamma_{N_{i}}}{H_{i,\rm eq}}=\frac{3\sqrt{3\lambda_{\phi}/2}}{8\pi{\rm B}_{i}^{2}}\left(\frac{m_{\nu_{i}}M_{\rm P}}{v_{h}^{2}}\right)>0.25\times\left(\frac{1/3}{{\rm B}_{i}}\right)^{2}\left(\frac{\phi_{0}}{10^{9}{\rm GeV}}\right)^{1/2}\left(\frac{m_{\nu_{i}}}{m_{\nu}}\right)$ (3.91)
@@
$\frac{m_{\nu_{1}}}{m_{\nu}}\lesssim 10^{-3}\left(\frac{\Delta N_{\rm eff}^{\rm obs}}{0.5}\right)^{3/2}\left(\frac{{\rm B}_{1}}{0.1}\right)^{2}\left(\frac{100}{\xi_{\phi}}\right)\ .$ (3.92)

### v3

@@ 3.3 Reheating after inflation
For $\phi$ well below the Planck scale, the scalar potential is approximated as $U\simeq{\tilde{V}}$. It can be expressed as a familiar form given in Eq. (2.16). For simplicity, we consider the case of $\lambda_{\phi}^{\prime}=\lambda_{\phi}$ (i.e., $\zeta_{h\phi}=0\ \Leftrightarrow\lambda_{\phi h_{i}}=0$) and $\lambda_{\phi h}=0$. Then, the mass of $\phi$-quanta for a given background value is given by
@@
For $\lambda_{\phi h_{i}}=\lambda_{\phi h}=0$, the production of the SM Higgs from either preheating processes or decays of $\phi$ particles is negligible. For sufficiently small $y_{N}$ which will be discussed in the subsequent discussion, the production of RHNs via preheating is also not efficient [60]. Hence, initially there can be population of $\phi$-particles and axi-Majorons ($a_{\phi}$) only  [^15] There are thermal particles with a temperature $T\sim H_{\rm e}/(2\pi)$ at the end of inflation [61]. We ignore this contribution which does not affect our argument. . As shown in Ref. [62], the self-preheating effect on $\phi$-particles is inefficient. The population of $a_{\phi}$ is only from the perturbative decay of $\phi$-particles through the kinetic term of $a_{\phi}$. Moreover, if $\Phi$ has nonzero angular momentum (i.e., a phase rotation), the preheating effect will be suppressed since $\phi$ does not pass through the origin during its oscillations and the violation of adiabaticity which is necessary for an efficient preheating becomes weak as a result. However, there might be a possibility of efficient energy transfer via axi-Majoron’s anomalous couplings to non-Abelian gauge fields due to large velocity of axi-Majoron after inflation [63]. Since the velocity of axi-Majoron is periodically changing as shown in Fig. 3, in regard of the polarization state of gauge field which is amplified the situation is similar to the case of parametric resonance in a conformal scalar theory [62]. The details will be investigated in other work. In any case, as long as the eventual decay of $\phi$-particle is the dominant source of the energy of the universe and the interaction of $\phi$ to RHNs is small enough to avoid the possibility of thermal-trapping, our argument is barely affected. Hence, in this work we ignore the effect of preheating.
In our scenario, $U(1)_{\rm PQ}$ symmetry is broken only in the gravity sector, and its effect is suppressed by $\alpha\xi_{\phi}\phi^{2}/M_{\rm P}^{2}$ and turned off as $\phi\ll c_{\phi}\chi$ [see Eq. (2.7)]. As a result, $a_{\phi}$ is lighter than $\phi$ for $\phi\lesssim M_{\rm P}$ with $\alpha\xi_{\phi}\ll 1$, and $\phi$ can always decay to $a_{\phi}$s via the kinetic term of $a_{\phi}$. The partial decay rate of the channel is given by
@@
From the seesaw formula for the mass parameters of the left-handed neutrinos [64, 65]  [^16] At low energy below the electroweak scale, $y_{\nu}$ should be replaced to the one obtained from a biunitarity transformation associated with physical mass-eigenstates. We ignore such details here, since it does not make changes of the main point of our argument. In this reason, we assumed $y_{\nu}$ is flavor-diagonal for simplicity.,
$m_{\nu_{i}}\approx\frac{y_{\nu_{i}}^{2}v_{u}^{2}}{2m_{N_{i}}}\ ,$ (3.64)
where $v_{u}\equiv h_{2,0}=v_{h}\sin\beta$, i.e., the VEV of the up-type Higgs. Then the constraint in Eq. (3.63) is translated as a constraint on $y_{\nu_{i}}$, which can be expressed as
$y_{\nu_{i}}^{2}$$\approx$ $2\sqrt{b_{i}\lambda_{\phi}}\left(\frac{m_{\nu_{i}}\phi_{0}}{v_{u}^{2}}\right)\ ,$ (3.65)
@@
For $\phi\gtrsim\phi_{\times}$ the universe is driven by a radiationlike energy component. In this case, the decay can happen effectively once $H\approx\Gamma^{\phi}/2$ is satisfied. If the subdominant angular motion is ignored, the expansion rate after inflation can be expressed such as
@@
$\frac{\phi_{*}}{\phi_{\times}}\approx\frac{9\lambda_{\phi}}{32\sqrt{2}\pi}\frac{M_{\rm P}}{\phi_{0}}\simeq\left(\frac{\xi_{\phi,0}}{10^{2}}\right)^{2}\left(\frac{\phi_{0}^{\rm ref}}{\phi_{0}}\right)\ .$ (3.70)
@@
where ${\tilde{\phi}}$ is understood as the amplitude of oscillation. In this regime, the universe is in a matter-domination era driven by nonrelativistic massive $\phi$-particles [^17] There is no kination era driven by the kinetic anergy caused by angular motion of $\Phi$. This is because the phase velocity of $\Phi$, i.e., $\dot{\theta}$ scales as $a^{-1}$ for $\phi\gtrsim\phi_{\times}$ and $a^{-3}$ for $\phi\lesssim\phi_{\times}$ in our scenario. Its energy contribution is always subdominant relative to the contribution from the radial mode.. Hence, requiring $H=(2/3)\Gamma^{\phi}$, one finds the oscillation amplitude ${\tilde{\phi}}_{*}$ at the epoch of the decay of $\phi$ as
${\tilde{\phi}}_{*}\approx\frac{\sqrt{6}\lambda_{\phi}}{24\pi}M_{\rm P}\simeq 4.7\times 10^{7}{\rm GeV}\times\xi_{\phi,0}^{2}\ ,$ (3.73)
@@
${\tilde{H}}_{*}\approx\frac{\sqrt{2}\lambda_{\phi}^{3/2}}{24\pi}\phi_{0}\simeq 3.5\times 10^{-4}{\rm GeV}\times\xi_{\phi,0}^{3}\left(\frac{\phi_{0}}{\phi_{0}^{\rm ref}}\right)\ .$ (3.74)
@@
$\xi_{\phi,0}\lesssim 680\left(\frac{0.1}{b_{N}}\right)^{1/5}\left(\frac{\phi_{0}}{\phi_{0}^{\rm ref}}\right)^{2/5}$ (3.77)
where we took $b_{N}\equiv b_{3}\approx b_{2}\gg b_{1}$. In the case of $\phi_{*}<\phi_{\times}$, the constraint on $\xi_{\phi,0}$ is much weaker since the would-be temperature of RHNs around the epoch of $\phi=\phi_{\times}$ is much lower. Hence, conservatively speaking, as long as Eq. (3.77) is satisfied, the PQ-field can avoid thermal-trap around the origin, and the PQ-symmetry is never restored during and after inflation.
@@
The details of the postinflation cosmology depends on whether $\phi_{*}$ is larger than $\phi_{\times}$ or not. So, we discuss those cases separately.
@@
$\xi_{\phi,0}<100\times\left(\frac{\phi_{0}}{\phi_{0}^{\rm ref}}\right)^{1/2}$ (3.78)
Once $\phi$ decays at the epoch of ${\tilde{H}}_{*}$, the universe will be filled with axi-Majorons and RHNs as a comparable or subdominant component. Depending on interaction rates, they may or may not reach thermal or kinetic equilibrium among themselves before the standard thermal bath is established from decays of RHNs. Initially, the production of $\phi$ via inverse decay is kinematically forbidden. Hence, thermalization before the decay of RHNs would be driven by interactions of axi-Majoron kinetic term, $y_{N_{i}}$, and $y_{\nu_{i}}$. The $\phi$-mediating self-interaction cross-section of axi-Majorons is given by
@@
It implies that $E_{a,\rm k}>{\tilde{E}}_{a,*}$ if $\xi_{\phi,0}\lesssim 100\left(\phi_{0}/\phi_{0}^{\rm ref}\right)^{1/2}$. Hence, for $\xi_{\phi,0}$ satisfying Eq. (3.78), axi-Majorons are never thermalized among themselves after their production in the decay of $\phi$.
@@
where $N=\left(\nu_{R},\nu_{R}^{c}\right)^{T}$ is a four-component Majorana field with the flavor index suppressed, and only relevant leading order terms are shown. When the condition in Eq. (3.63) is satisfied for the RHN-channels of the decay of the inflaton, the cross sections of the various interaction between the axi-Majorons and the RHNs are smaller than the one for axi-Majoron self-interactions at high energy limit($s\approx m_{\phi,0}$) at least by a factor of about $b_{i}/3$ (see Appendix C). Moreover, the number density of RHNs is also suppressed by ${\rm B}_{i}$. As a result, for $\xi_{\phi,0}$ satisfying Eq. (3.78), the associated interaction rates cannot catch up the expansion rate for $H<H_{\times}\equiv H(\phi_{\times})$.
@@
$\frac{y_{\nu_{i}}^{4}}{y_{N_{i}}^{4}}=\left[\frac{m_{\nu}\phi_{0}}{v_{u}^{2}\sqrt{b_{i}\lambda_{\phi}}}\left(\frac{m_{\nu_{i}}}{m_{\nu}}\right)\right]^{2}\approx\frac{10^{3}}{b_{i}\xi_{\phi,0}^{2}\sin^{4}\beta}\left[\frac{\phi_{0}}{\phi_{0}^{\rm ref}}\frac{m_{\nu_{i}}}{m_{\nu}}\right]^{2}\ ,$ (3.84)
one may think that the interaction $y_{\nu_{i}}$ can cause much more efficient scatterings and annihilations of $N_{i}$s to SM particles. However, the decay rate of a RHN mass eigenstate $N_{i}$ is given by
$\Gamma_{N_{i}}\simeq\frac{y_{\nu_{i}}^{2}m_{N_{i}}}{8\pi}=\frac{m_{\nu_{i}}}{4\pi}\left(\frac{m_{N_{i}}}{v_{u}}\right)^{2}=\frac{b_{i}\lambda_{\phi}m_{\nu_{i}}}{4\pi}\left(\frac{\phi_{0}}{v_{u}}\right)^{2}\ ,$ (3.85)
where RHNs were assumed to be much heavier than the particles of the Higgs doublets. Also, for the heavy RHN mass-eigenstates which we assume to be $N_{2,3}$ we can take $m_{\nu_{i}}\approx m_{\nu}$ while both of the normal or inverted mass hierarchy of the left-handed neutrino are still allowed, depending on the details of $y_{\nu}$ [14]. Then, as long as $\tan\beta\gtrsim 1$ which may have to be the case to match electroweak precision data including the stability of the electroweak vacuum [67], for $\xi_{\phi,0}$ in the range we are interested in, with $\phi_{0}\sim\phi_{0}^{\rm ref}$, heavy RHN mass-eigenstates decay well before their self-interactions/annihilations involving either axi-Majorons or SM particles become relevant. The lightest RHNs can also be decoupled from SM particles before their decay if $m_{\nu_{1}}\ll m_{\nu}$ even in the case of $\xi_{\phi,0}\ll\mathcal{O}(10^{2})$. Therefore, once produced, axi-Majorons and RHNs are expected to evolve independently without thermalization at least until the epoch of RHNs decay as long as $m_{\nu_{1}}\ll m_{\nu_{2}}\approx m_{\nu_{3}}\approx m_{\nu}$.
If it were stable, a RHN mass-eigenstate becomes nonrelativistic when the expansion rate becomes
@@
$\frac{\Gamma_{N_{i}}}{{\tilde{H}}_{i,\rm non}}=\frac{3}{\sqrt{2\lambda_{\phi}}}\frac{m_{\nu_{i}}\phi_{0}}{v_{u}^{2}}\simeq\frac{66}{\xi_{\phi,0}\sin^{2}\beta}\left(\frac{m_{\nu_{i}}}{m_{\nu}}\right)\left(\frac{\phi_{0}}{\phi_{0}^{\rm ref}}\right)\ .$ (3.87)
@@
$\frac{\Gamma_{N_{i}}}{{\tilde{H}}_{i,\rm eq}}=\frac{3}{2\sqrt{\lambda_{\phi}}{\rm B}_{i}^{2}}\left(\frac{m_{\nu_{i}}\phi_{0}}{v_{u}^{2}}\right)\approx\frac{51}{{\rm B}_{i}^{2}\xi_{\phi,0}\sin^{2}\beta}\left(\frac{m_{\nu_{i}}}{m_{\nu}}\right)\left(\frac{\phi_{0}}{\phi_{0}^{\rm ref}}\right)\ .$ (3.90)
If one of RHN mass-eigenstates is responsible for eventual recovering the SM thermal bath, it would have to satisfy $\Gamma_{N_{i}}/{\tilde{H}}_{i,\rm eq}<0.1\times\left(\Delta N_{\rm eff}^{\rm obs}/0.5\right)^{3/2}$ with $\Delta N_{\rm eff}^{\rm obs}$ being the observational bound on the extra radiation (see Sec. 5). Combined with the condition in Eq. (3.78), the branching ratio ${\rm B}_{i}$ is then constrained as
${\rm B}_{i}\gtrsim\frac{0.40}{\sin\beta}\times\left(\frac{0.5}{\Delta N_{\rm eff}^{\rm obs}}\right)^{3/4}\left(\frac{m_{\nu_{i}}}{m_{\nu}}\right)^{1/2}\left(\frac{\phi_{0}}{10^{9}{\rm GeV}}\right)^{1/4}\ ,$ (3.91)
@@
$\frac{m_{\nu_{1}}}{m_{\nu}}<2\times 10^{-4}\times\sin^{2}\beta\left(\frac{\Delta N_{\rm eff}^{\rm obs}}{0.5}\right)^{3/2}\left(\frac{\xi_{\phi,0}}{10}\right)\left(\frac{{\rm B}_{1}}{0.1}\right)^{2}\left(\frac{\phi_{0}^{\rm ref}}{\phi_{0}}\right)\ .$ (3.92)
Then, it is necessary to have $\phi_{0}\sim\phi_{0}^{\rm ref}$ for the observed relic density of dark matter (see Sec. 5), and heavy states are expected to decay around or before the epoch they become nonrelativistic, and never dominate the universe.
If heavy RHN states decay before or around the epoch they become nonrelativistic, their decay rate is suppressed by the boosting effect. For simplicity, we may take ${\rm B}_{N}\equiv{\rm B}_{3}\approx{\rm B}_{2}$, $\Gamma_{N}\equiv\Gamma_{N_{3}}\approx\Gamma_{N_{2}}$, $m_{\nu_{N}}\equiv m_{\nu_{3}}\approx m_{\nu_{2}}$, $m_{N}\equiv m_{N_{3}}\approx m_{N_{2}}$. Then, the expansion rate at the epoch of the decays of heavy RHNs is estimated as
$H_{N}=\frac{f_{N}(\xi_{\phi,0},\phi_{0})}{2}\Gamma_{N}\ ,$ (3.93)
@@
$f_{N}(\xi_{\phi,0},\phi_{0})\equiv\left[\frac{{\tilde{H}}_{*}}{\Gamma_{N}/2}\left(\frac{m_{N}}{m_{\phi,0}}\right)^{2}\right]^{1/3}\simeq 0.3\xi_{\phi,0}^{1/3}\left(\frac{\phi_{0}^{\rm ref}}{\phi_{0}}\right)^{1/3}\ ,$ (3.94)
@@
If $\phi_{*}>\phi_{\times}$, i.e., $\xi_{\phi,0}>100\sqrt{\phi_{0}/\phi_{0}^{\rm ref}}$ (see Eq. (3.70)), $\Phi$ field decays before falling into the region of the true vacuum position. Even in this case, $N_{i}$ can be kept out of kinetic or thermal equilibrium as long as $\xi_{\phi,0}\sim\mathcal{O}(10^{2})$ with $\phi_{0}\approx\phi_{0}^{\rm ref}$. The expansion rate at the epoch when $N_{i}$ becomes nonrelativistic is found as
@@
$\frac{\Gamma_{N_{i}}}{H_{i,\rm eq}}=\frac{3\sqrt{3\lambda_{\phi}/2}}{8\pi{\rm B}_{i}^{2}}\left(\frac{m_{\nu_{i}}M_{\rm P}}{v_{u}^{2}}\right)>\frac{0.25}{\sin^{2}\beta}\times\left(\frac{1/3}{{\rm B}_{i}}\right)^{2}\left(\frac{\phi_{0}}{10^{9}{\rm GeV}}\right)^{1/2}\left(\frac{m_{\nu_{i}}}{m_{\nu}}\right)$ (3.98)
@@
$\frac{m_{\nu_{1}}}{m_{\nu}}\lesssim 10^{-3}\times\sin^{2}\beta\left(\frac{\Delta N_{\rm eff}^{\rm obs}}{0.5}\right)^{3/2}\left(\frac{{\rm B}_{1}}{0.1}\right)^{2}\left(\frac{100}{\xi_{\phi,0}}\right)\ .$ (3.99)